### v1

# Multi-Modal Federated Learning for Cancer Staging over Non-IID Datasets with Unbalanced Modalities

Abstract
The use of machine learning (ML) for cancer staging through medical image analysis has gained substantial interest across medical disciplines. When accompanied by the innovative federated learning (FL) framework, ML techniques can further overcome privacy concerns related to patient data exposure. Given the frequent presence of diverse data modalities within patient records, leveraging FL in a multi-modal learning framework holds considerable promise for cancer staging. However, existing works on multi-modal FL often presume that all data-collecting institutions have access to all data modalities. This oversimplified approach neglects institutions that have access to only a portion of data modalities within the system. In this work, we introduce a novel FL architecture designed to accommodate not only the heterogeneity of data samples, but also the inherent heterogeneity/non-uniformity of data modalities across institutions. We shed light on the challenges associated with varying convergence speeds observed across different data modalities within our FL system. Subsequently, we propose a solution to tackle these challenges by devising a distributed gradient blending and proximity-aware client weighting strategy tailored for multi-modal FL. To show the superiority of our method, we conduct experiments using The Cancer Genome Atlas program (TCGA) datalake considering different cancer types and three modalities of data: mRNA sequences, histopathological image data, and clinical information.
Index Terms:
Multi-modal federated learning, gradient blending, client weighting, heterogeneous modalities, non-iid data
I Introduction
Leveraging medical images to derive machine learning (ML) models for aiding in cancer staging and developing effective treatment strategies has shown significant success [1]. Traditionally, despite the existence of multiple modalities of data (e.g., computed tomography (CT) scans [2, 3] and histopathological images [4]), ML models (e.g., MLPs, CNNs, and LSTMs) are trained with one of data modalities [5, 6, 7]. However, as research revealed the benefits of multi-modal ML [8], modern approaches have utilized multiple data types/modalities such as mRNA sequences and clinical information together to obtain ML models for medical applications [9].
The majority of existing studies on multi-modal ML within the medical domain [10, 11, 12] presume a centralized ML model training architecture: ML models are trained at a module (e.g., a server) with access to a repository of patient data. However, since patient data is often distributed across various data collection units (e.g., hospitals, clinics), the centralized approach necessitates the transfer of patient data across the network to a centralized location, raising data privacy concerns. Researchers have thus explored alternative ML approaches, a notable example of which is federated learning (FL) [13, 14].
In FL, each data collecting institution, referred to as a client, independently trains a local ML model using its data. These local ML models are periodically aggregated by a server to construct a global model. The global model is disseminated to clients to commence their next local ML model training cycle. Consequently, FL replaces the transmission of raw data across the network with the transmission of ML models, thus preserving data privacy. As a result, the adoption of FL within the medical domain is gaining substantial traction [15, 16, 17, 18, 19]. Given the diversity of data types/modalities across FL clients in medical domain, such as hospitals and clinics, the prospect of employing FL within a multi-modal learning context is particularly enticing. Although FL and multi-modal ML have been jointly studied recently [17, 20], a common constraint in these studies is the assumption that each client has access to all data modalities. In real-world scenarios, this assumption often does not hold: for example, a hospital may collect mRNA sequences and histopathological images, while a clinic may collect only CT scans and clinical information.
In this work, we take an initial step towards implementing multi-modal FL in a scenario where clients not only have non-iid (non-independent and identically distributed) data, but also possess an uneven number of data modalities. We present a system model that eliminates the need for clients to possess identical sets of data modalities by employing a versatile distributed encoder-decoder architecture. Within this framework, the encoders are tasked with the extraction of features from each distinct data modality at each client. These encoders and decoders/classifiers will undergo separate aggregations in order to address the heterogeneity of modalities across clients. This approach ensures that all trainable components relevant to a specific data modality are collectively trained using the data of that particular modality across all clients. The final global model resulting from this training can then be applied over the combination of all data modalities.
We show that while our proposed system model holds considerable promise, it confronts a series of challenges. A notable concern is the disparity in the training and convergence speeds of various data modalities among different clients. This variability can lead to certain encoders rapidly exhibiting a biased behavior towards local data of clients, which can impede the global model’s convergence. [^1] This phenomenon, in turn, may result in overfitting toward specific data labels, particularly in the context of generic FL classification tasks, due to the imbalanced effect of each modality during local ML training. To tackle this challenge, we introduce a novel approach termed distributed gradient blending (DGB) for FL. Although conceptually distinct, this approach draws inspiration from a seminal work on centralized multi-model ML [21]. Our proposed DGB method capitalizes on information pertaining to the performance of participating clients in FL and the characteristics of their respective modalities to determine effective back-propagation weights for local training of each encoder at each client. To further enhance the robustness of our proposed method against non-iid data distribution across the clients, we propose proximity-aware client weighting (PCW). PCW considers the quality of data at clients to further enhance the parameters utilized in DGB.
The major contributions of this work are summarized below:
  • We present a novel system model that accommodates an uneven number of data modalities across clients in FL. This introduces an additional layer of data heterogeneity, superimposed upon the inherent non-iid nature of per-modality data distribution across the clients in FL. The introduction of this system model allows real-world applications of multi-modal FL, particularly within the domain of healthcare.
  • We identify challenges associated with achieving high-performance ML models within our system. To tackle a crucial challenge related to disparate convergence rates across data modalities of various clients, we introduce DGB. This approach is specifically designed for FL systems with clients having uneven numbers of data modalities. DGB harnesses system-wide information, namely, the training and validation losses, and data characteristics, to dynamically adjust the learning-rate of each modality at each FL client.
  • To enhance the convergence speed and thus improving the communication efficiency of our approach, we introduce PCW. PCW regulates the calculation of DGB parameters with respect to the varying qualities of data present across the FL clients (i.e., the fidelity of estimate).
  • We conduct numerical evaluations on a dataset of three cohorts gathered by The Cancer Genome Atlas program (TCGA - https://www.cancer.gov/tcga) for three different cancer types (breast, lung, and liver) and demonstrate that our method can lead to significant performance gains as compared to various baseline methods in multi-modal FL.
II Related Work
In this section, we review the existing literature on cancer staging using ML methods and the expansions and contributions of FL to this active research area. Henceforth, we use the terminology of ML models to refer to deep neural networks.
Single-Modal Cancer Staging. In most of the early works [5, 6, 7, 3, 2, 4], only one modality of data is used in order to train the ML models. Thus, the ML models developed in these works take an input datapoint of a specific modality (e.g., PET/CT image or histopathological image) and pass it through – either raw or preprocessed – to predict its class (e.g., the cancer stage). As a result, most of the recent efforts in this area  [6, 7], have been focused on optimizing the ML model architecture and developing effective data preprocessing methodologies. However, the aformentioned studies overlook the proliferation of diverse data modalities that hold the potential to significantly enhance the accuracy of ML models for cancer prediction and staging. Moreover, these studies often make the assumption of unrestricted access to a comprehensive reservoir of patient data, the acquisition of which is becoming increasingly challenging due to the concerns regarding data privacy. In this work, we aim to address both of these issues by proposing multi-modal FL and resolving the challenges it comes with.
Multi-Modal Cancer Staging. Researchers have investigated the advantages of leveraging diverse data modalities for cancer prognosis [22, 23, 9, 24, 25]. For instance, in [9], the authors demonstrated the significant impact of incorporating histopathological images, gene expression data, and clinical data on the performance of ML models used for cancer staging. Also, the utilization of CT scan data, clinical data, and selected deep-learning-based radiomics data in cancer recurrence prediction has yielded similar ML prediction enhancements in [24]. Nevertheless, these studies mostly operate within a centralized ML training architecture, where the ML training module, often hosted on a central server, is assumed to have access to a large-scale repository of patient data. This assumption limits the applicability of these findings in real-world scenarios where patient data is distributed across multiple medical institutions, pooling of which in a centralized location raises major concerns regarding leakage of patients’ information. Our contribution to this literature lies in the proposition of a distributed ML architecture inspired by FL that eliminates the need for patients data transfer, while explicitly accommodating the diversity of data modalities across distinct data collection institutions.
Federated Learning. Motivated by the privacy concerns associated with the transfer of data across the network, there has been a notable interest in the application of distributed ML techniques, with particular attention directed towards FL, in various fields of studies [26, 27]. Nevertheless, majority of works on FL [28, 29] presume the existence of single modality of data across the clients, and thus cannot be easily deployed in systems with multiple modalities of data.
Multi-modal FL for Cancer Classification. Multi-modal FL has recently attracted notable attention within the research community of various fields, such as healthcare and machine vision [17, 30, 31, 32, 20, 33], among which, considering medical applications, several endeavors stand out [17, 33]. In [17], an FL system effectively incorporates both X-ray and Ultrasound modalities. In [33], multi-modal FL is proposed with skin lesion images and their corresponding clinical data for melanoma detection. However, a common thread runs through all of the previous works, implying both to the general context of multi-modal FL and the cases where it has been applied to medical topics: they all operate under the assumption that every client has access to (and collects) all of data modalities. In practice this assumption often encounters limitations in the medical domain as data collection institutions often gather only a subset of data modalities; for instance, one hospital may gather different gene expression data and clinical data, while another may collect CT scans. Our work is among the first in the literature to introduce an effective FL paradigm across clients with an uneven distribution/number of data modalities.
III Network and Learning Model
In this section, we first present our network model of interest (Sec. III-A). Subsequently, we offer a concise discussion on conventional uni-modal FL (Sec. III-B). Our contribution unfolds as we extend the conventional FL paradigm by introducing a novel FL system designed to accommodate a network of clients that possess an uneven distribution/number of data modalities (Sec. III-C). We conclude this section by providing an overview of the pivotal challenges that necessitate addressing to achieve high-performance ML models within the context of our proposed FL system (Sec. III-D).
III-A Network Model
We consider a network of $N$ institutions/clients gathered via the set $\mathcal{N}=\{1,\cdots,N\}$, $|\mathcal{N}|=N$. Each institution $n\in\mathcal{N}$ is assumed to possess a dataset $\mathcal{D}_{n}=\mathcal{D}_{n}^{\mathsf{Tr}}\cup\mathcal{D}_{n}^{\mathsf{Va}}$ with size $D_{n}\triangleq|\mathcal{D}_{n}|$ where $\mathcal{D}_{n}^{\mathsf{Tr}}$ is the dataset used for training and $\mathcal{D}_{n}^{\mathsf{Va}}$ is the dataset used for validation. We also denote the union of datasets collected across the institutions, called global dataset, as $\mathcal{D}\triangleq\cup_{n\in\mathcal{N}}{\mathcal{D}_{n}}$ with $D\triangleq|\mathcal{D}|$ datapoints. [^2] Without loss of generality, as commonly presumed in the literature of FL [32, 20], we assume that datasets across the institutions are non-overlapping.
III-B Conventional Uni-modal Federated Learning
In the conventional uni-modal FL framework, the datasets held by institutions, and consequently the global dataset, consist solely of a single data modality (e.g., image data) [13, 26, 27]. FL conducts ML model training through a series of local (model) training rounds indexed by $k\in\in\{0,\cdots,K-1\}$, each leading to a global (model) aggregation step, indexed by $t\in\{0,\cdots,T-1\}$. Within this context, each institution $n$ trains a local ML model, represented by a parameter vector $\bm{\omega}_{n}$. The initial local training round commences with the server initializing a global model $\bm{\omega}^{(t)}$ at $t=0$ and disseminating it among the institutions. Subsequently, each institution $n$ synchronizes its local model as $\bm{\omega}_{n}\leftarrow\bm{\omega}^{(t)}$ for $t=0$ and utilizes it as the initial model for local training, i.e., $\bm{\omega}^{(t),0}_{n}\leftarrow\bm{\omega}_{n}$. Following this initialization, each client undergoes $K$ local stochastic gradient descent (SGD) iterations, utilizing its own dataset to obtain $\bm{\omega}^{(t),K}_{n}$. The interim ML updates, for $10\leq k\leq K-$ 1, are computed as $\bm{\omega}^{(t),k+1}_{n}=\bm{\omega}^{(t),k}_{n}-\eta^{(t),k}\widetilde{\nabla}\mathcal{L}_{n}(\bm{\omega}^{(t),k}_{n})$, where $\eta^{(t),k}$ denotes the step-size/learning-rate, and $\mathcal{L}_{n}(\cdot)$ denotes the local loss function of the institution $n$, which quantifies the predictive accuracy of the ML model over its local dataset. Further, ~$\widetilde{\nabla}\mathcal{L}_{n}(\cdot)$ denotes the stochastic gradient obtained by sampling a random mini-batch of data from the local dataset $\mathcal{D}^{\mathsf{Tr}}_{n}$.
Upon acquiring $\bm{\omega}^{(t),K}_{n}$ at each institution $n$, the institution transmits it to the server through upstream/uplink communications. Subsequently, the server aggregates the gathered ML models of institutions $\{\bm{\omega}^{(t),K}_{n}\}_{n\in\mathcal{N}}$ to form a new global model $\bm{\omega}^{(t+1)}$ via the following process:
$\bm{\omega}^{(t+1)}=\sum_{n\in\mathcal{N}}{D_{n}\bm{\omega}_{n}^{(t),K}}\big{/}{D}.$ (1)
This global model is broadcast to institution, using which they synchronize their local models as mentioned above and initiate the next round of local ML training and global aggregation.
III-C Federated Learning over Unbalanced Modalities
III-C1 Multi-modal FL under Unbalanced Modalities
In real-world scenarios, institutions may have access to various subsets of modalities. This phenomenon makes the conventional multi-modal FL, which is based on naïve extension of uni-modal FL to multi-modal settings [31, 32, 20] impractical, as these works presume the existence of the same set of data modalities across all the institutions. In particular, in a network of institutions with non-uniform number of data modalities, the number of encoders used across them will differ. Further, the size of the input feature fed to the classifier will also vary across institutions. These variabilities make the naïve ML training and aggregation methods of [31, 32, 20] ineffective. Consequently, we propose a generalized distributed encoder-decoder architecture, which increases the flexibility of model aggregations at the server by aggregating (i) each modality’s encoder and (ii) classifiers of different existing modality combinations across the institutions separately. A schematic of our proposed system and model aggregation strategy is depicted in Fig. 1(a) and 1(b), respectively.
Fig. 1: (a) Our system model: each institution $n$ can have access to a subset of data modalities ($\mathcal{M}_{n}\subseteq\mathcal{M}$). (b) The server gathers encoders and classifiers and aggregates them to global models and sends them back to institutions.
Let $\mathcal{M}$ and $\mathcal{Y}$ denote the sets encompassing all data modalities within our network distributed across the institutions, and the data labels (i.e., cancer stages in our scenario), respectively. Moreover, let $\mathcal{M}_{n}\subseteq\mathcal{M}$ denote the subset of modalities held by institution $n$, which are present in its local dataset $\mathcal{D}_{n}$. Considering each institution $n$, each data point $d\triangleq(\bm{X},\bm{y})\in\mathcal{D}_{n}$ is associated with a feature vector $\bm{X}$ and a label $\bm{y}$, where $\bm{X}=\{\bm{x}_{m}\}_{m\in\mathcal{M}_{n}}$ contains the set of features corresponding to each modality $m$ possessed by the institution. Additionally, $\bm{y}\in\mathcal{Y}$ is represented in a one-hot encoded format, capturing the cancer stage. As will be formalized next, to handle clients with an uneven distribution of modalities, we will put a greater emphasis on improving the local training of encoders. This approach diverges from the recent methodologies in multi-modal FL [32], where the emphasis lies on aggregating the classification layers – and co-attention layers – which adapt well to scenarios where clients have access to all the modalities. This paradigm shift allows our approach to extend the distributed ML training framework of FL for each modality to all institutions possessing data pertaining to that specific modality. In particular, in our approach, the encoder used for each modality would benefit from the commonality of data across the institutions possessing that modality. This, in turn, eliminates the restricting assumption that institutions must possess all data modalities.
Considering encoder parameters/models $\{\bm{\omega}_{m,n}\}_{m\in\mathcal{M}_{n}}$ and the classifier parameters $\bm{\varpi}_{n}$ at institution $n$ (see Fig. 2), the ultimate predicted label $\widehat{y}$ for a datapoint $d\in\mathcal{D}_{n}$ is given by
$\widehat{y}=\bm{\varpi}_{n}\Big{(}\big{[}\bm{\omega}_{m_{1},n}(\mathbf{x}_{m_{1}}):\bm{\omega}_{m_{2},n}(\mathbf{x}_{m_{2}}):\cdots:\bm{\omega}_{|\mathcal{M}_{n}|,n}(\mathbf{x}_{m_{|\mathcal{M}_{n}|}})\big{]}\Big{)},$
where $\bm{\omega}_{m_{i},n}(\mathbf{x}_{m_{i}})$, for each $i\in\{1,\cdots,|\mathcal{M}_{n}|\}$, represent the encoder output upon being fed with data feature of modality $m_{i}$. Also, the function $\bm{\varpi}_{n}(\cdot)$ encompasses the description of the decoder/classifier, which takes as its input the concatenated/fused feature of all the modalities available at the institution and produces a predicted label as its output.
Fig. 2: A schematic of the encoder-decoder structure at an arbitrary institution $n$ with three data modalities.
III-C2 Machine Learning Task Formulation
We consider the loss function at institution $n$ for an arbitrary local model parameter realization $\bm{\omega}_{n}$ as $\mathcal{L}_{n}(\bm{\omega}_{n};\mathcal{D}_{n})=\frac{1}{|\mathcal{D}_{n}|}\sum_{d\in\mathcal{D}_{n}}\ell(\bm{\omega}_{n};d)$ where $\ell(\bm{\omega}_{n};d)$ is the ML model loss function evaluated on datapoint $d$. In this work, we use cross-entropy to measure the loss [34]. Also, we use $\bm{\omega}_{n}\triangleq\left[\{\bm{\omega}_{m,n}\}_{m\in\mathcal{M}_{n}}:\bm{\varpi}_{n}\right]$ to denote the concatenation of the encoders model parameters in conjunction with the decoder for institution $n$.
Since in our system of interest, institutions have different combinations of modalities, to capture the present combinations of modalities in the system, we introduce a set [^3] Note that $\mathcal{P}(\mathcal{M})$ is a set and thus only has non-identical elements.
$\mathcal{P}(\mathcal{M})\triangleq\{\mathcal{M}_{n}\}_{n\in\mathcal{N}}.$ (2)
If all the institutions have the same modality ($\mathcal{M}_{n}=\mathcal{M},\forall n$), then $\mathcal{P}(\mathcal{M})=\mathcal{M}$. Subsequently, since institutions with the same combinations of modalities use the same architecture of classifiers, we denote $\widehat{\bm{\varpi}}_{C}$ as the common architecture of the classifier used in the system, for all institutions $n$ with modality combination $C\in\mathcal{P}(\mathcal{M})$, for which $|\widehat{\bm{\varpi}}_{C}|=|\bm{\varpi}_{n}|$, $\forall n$, where |.||.|| . | captures the number of classifier parameters. [^4] For example, if the system has one institution with modality $m_{1}$, one institution with modality $m_{2}$, one institution with modality $m_{3}$, one institution with modality $m_{1},m_{2}$, and one institution with modality $m_{1},m_{2},m_{3}$, then, $\mathcal{P}(\mathcal{M})=\{\{m_{1}\},\{m_{2}\},\{m_{3}\},\{m_{1},m_{2}\},\{m_{1},m_{2},m_{3}\}\}$. In this case, $C$ can be any of the five sets that constitute $\mathcal{P}(\mathcal{M})$.
Subsequently, we can define the global loss function as
$\mathcal{L}(\bm{\omega})\triangleq\sum_{n\in\mathcal{N}}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\mathcal{L}_{n}(\bm{\omega}_{n}).$ (3)
Therefore, our goal is to find a set of global model parameters, consisting of a set of model parameters for encoders used for different modalities and the decoders. Mathematically, we are interested in obtaining $\bm{\omega}^{\star}$, which is given by
$\bm{\omega}^{\star}=\big{\{}\{\widehat{\bm{\omega}}^{\star}_{m}\}_{m\in\mathcal{M}},\{\widehat{\bm{\varpi}}_{C}^{\star}\}_{C\in\mathcal{P}(\mathcal{M})}\big{\}}\triangleq\operatorname*{arg\,min}_{\bm{\omega}\in\mathbb{R}^{|\bm{\omega}|}}\mathcal{L}(\bm{\omega}),$ (4)
where $|\bm{\omega}|=\sum_{m\in\mathcal{M}}|\widehat{\bm{\omega}}_{m}|+\sum_{C\in\mathcal{P}(\mathcal{M})}|\widehat{\bm{\varpi}}_{C}|$.
III-C3 Local ML Model Training
At each institution $n$, the local model training can be described via a sequence of mini-batch SGD iterations. In particular, given a local model at an institution at SGD iteration $k$, at global model aggregation step $t$ (i.e., $\bm{\omega}_{n}^{(t),k}$), the next local model is obtained as follows:
[…]
where
~$\widetilde{\nabla}\mathcal{L}_{n}(\bm{\omega}^{(t),k}_{n};\mathcal{B}_{n}^{(t),k})\triangleq\frac{1}{|\mathcal{B}_{n}^{(t),k}|}\sum_{d\in\mathcal{B}_{n}^{(t),k}}\nabla\ell(\bm{\omega}^{(t),k}_{n};d),$ (6)
and $\bm{\omega}_{n}^{(t),k}=\big{\{}\{\bm{\omega}_{m,n}^{(t),k}\}_{m\in\mathcal{M}_{n}}:\bm{\varpi}^{(t),k}_{n}\big{\}}.$ In (6), $\mathcal{B}_{n}^{(t),k}$ denotes a mini-batch of data sampled uniformly at random from $\mathcal{D}_{n}^{\mathsf{Tr}}$ and $\eta^{(t),k}$ is the SGD learning-rate. Assuming that institutions conduct $K$ SGD iterations before transmitting their models to the server, we let $\bm{\omega}_{n}^{(t),K}=\big{\{}\{\bm{\omega}_{m,n}^{(t),K}\}_{m\in\mathcal{M}_{n}}:\bm{\varpi}^{(t),K}_{n}\big{\}}$ denote the final local encoder and decoder parameters at institution $n$.
III-C4 Local Model Transfer and Global Model Aggregations
At the end of each local training round, after receiving the parameters of the encoders and decoders from the institutions, the server then aggregates them as follows to obtain a unique encoder parameter set for each modality and a unique decoder parameter set for each of the existing combinations of modalities in the system:
$\widehat{\bm{\omega}}_{m}^{(t+1)}$ $=\frac{\sum_{n=0}^{N}a_{m,n}|\mathcal{D}_{n}|\bm{\omega}_{m,n}^{(t),K}}{|\mathcal{D}|},\forall m\in\mathcal{M},$ (7)
$\widehat{\bm{\varpi}}_{C}^{(t+1)}$ $=\frac{\sum_{n=0}^{N}a_{C,n}|\mathcal{D}_{n}|\bm{\varpi}_{n}^{(t),K}}{|\mathcal{D}|},\forall C\in P(\mathcal{M}),$ (8)
where $a_{m,n}$, $\forall n,m$ is a binary indicator which takes the value of $1$ if institution $n$ has access to modality $m$; otherwise $a_{m,n}=0$. Similarly, $a_{C,n}$, $\forall n,C$, is a binary indicator, which is $1$ if institution $n$ has modality combination $C=\{m_{i}\}_{i\in\{1,\cdots,\mathcal{M}_{n}\}}$; otherwise $a_{C,n}=0$. After conducting the model aggregation step, the new sets of encoders and decoders are sent back in the downstream link to the institutions, which are then used to synchronize the local models. In particular, each institution $n$ with modality combination $C$ will conduct ${\small\bm{\omega}_{m,n}^{(t+1),0}\leftarrow\widehat{\bm{\omega}}_{m}^{(t+1)}}$ and ${\small\bm{\varpi}_{n}^{(t+1),0}\leftarrow\widehat{\bm{\varpi}}_{C}^{(t+1)}}$ (i.e., $\bm{\omega}_{n}^{(t+1),0}=\big{[}\{\widehat{\bm{\omega}}_{m}^{(t+1)}\}_{m\in\mathcal{M}_{n}}:\widehat{\bm{\varpi}}_{C}^{(t+1)}\big{]}$). This synchronization will be followed by SGD iterations described in (5) and then the next global aggregation step. This procedure continues until reaching a desired accuracy.
Fig. 3: Loss plots of the global models based on the modalities possessed across institutions in the first 100 global aggregation rounds. Plots consider various modality combinations: (a) image and mRNA modalities, (b) clinical and mRNA, and (c) image and clinical. The shaded areas represent the noise (20% standard deviation) for each combination of modalities tested. A key observation is the occasional degradation of the performance of the multi-modal ML models compared to their single-modal counterparts. This observation is consistent with the claims in [31] which emphasises on the necessity of syncing the convergence rates of the modalities in a multi-modal FL scenario. Further, our method leads to the balance of convergence rates across the modalities.
III-D Key Challenges and Design Criteria
Our proposed system model contains nuances in decoupling the aggregation of encoders and decoders, which is necessary in handling institutions with an unbalanced number of modalities. However, achieving a high prediction performance across this system faces a major obstacle caused by straggling modalities. We demonstrate this phenomenon in Fig. 3, in which we conduct ML model training according to Sec. III-C, which implies training all institutions with unified learning-rates (labelled as Naïve FL in the plots). We consider a dataset of three modalities for cancer staging, consisting of mRNA sequences, histopathological images, and clinical information and presume the existence of $9$ institutions, where $3$ have access to only one modality of data, $3$ to two modalities of data, and $3$ to all modalities of data. The details of the dataset used for patients are given in Sec. V-A. As can be seen in the figure, the performance of some modalities is notably better than the others, as the losses are decreasing faster for those modalities but the same does not apply to others. Furthermore, during the local training phase, the ML model of some modality combinations has converged notably faster than the rest. Also, some modalities struggle to improve their loss and accuracy during the ML model training. We next aim to propose a framework to address these issues, the result of which is labelled as ours in the plots in Fig. 3. It can be observed from the curves that our method balances the convergence rates of various modality combinations, leading to a notable performance gain as verified later in Sec. V.
IV Distributed Gradient Blending and Proximity-Aware Client Weighting Approach
In the following, we first provide an overview of our methodology to address the above-described issues of multi-modal FL (Sec. IV-A). We then describe the two key components of our method which consists of distributed gradient blending (Sec. IV-B) and proximity-aware client weighting (Sec. IV-C).
IV-A Tackling Multi-Modal FL with Unbalanced Modalities
Considering the discussion in Sec. III-D, we aim to propose a methodology for multi-modal FL that alters the local gradient computations to solve the problem of different convergence rates of various modalities. We first introduce the two key components of our method:
  • Distributed Gradient Blending (DGB): To remove the barrier of convergence speeds, DGB weights the gradients that are being back-propagated through each encoder in an institution’s model to unify the convergence rate of different modalities. DGB achieves this through acquiring a set of train and validation losses from each institution at the end of each local ML training round, which is in turn used to calculate two metrics of overfitting and generalization. These two metrics are used to obtain the aforementioned weights for the gradients, which can accelerate or slacken the ML training rate of different modalities at various institutions.
  • Proximity-Aware Client Weighting (PCW): PCW mitigates the performance degradation of DGB caused by the heterogeneity of data across institutions, which leads to the bias of the locally trained ML models, especially as the length of local ML model training (i.e., $K$) grows. PCW tunes the weights/importance of ML models of institutions during the system-wide DGB weight calculation based on the trajectory of models’ –both local and global– gradients. Specifically, PCW assigns greater importance to institutions whose local ML parameters’ trajectory exhibit a closer alignment with that of the global model. This signifies that institutions with superior data quality (i.e., diverse local datasets that better encapsulate the global data distribution, and thus do not induce bias onto the locally trained ML model), will have a more pronounced influence on the calculation of DGB weights. By mitigating the impact of local ML model bias, PCW makes DGB robust against the increase in the number of local SGD iterations $K$ in each global aggregation (i.e., reductions in uplink transmission frequency).
We next describe the working mechanism of DGB and PCW.
IV-B Distributed Gradient Blending (DGB)
To realize DGB, we first adapt and extend the notions of generalization and overfitting, defined in [21], to the federated setting. These notions provide insights into the effect of local ML training with a given unified step size across modalities on the performance of the local ML models. In DGB, we use both the validation loss and training loss across institutions to compute the local gradient back-propagation weights for the upcoming local training rounds. In particular, let $\mathcal{L}_{n}^{\mathsf{Va}}$ denote the validation loss for institution $n$’s model at global aggregation step $t$ at SGD iteration $k$, which is
$\mathcal{L}_{n}^{\mathsf{Va}}(\bm{\omega}_{n}^{(t),k})\triangleq$ $\mathcal{L}_{n}(\bm{\omega}_{n}^{(t),k};\mathcal{D}^{\mathsf{Va}}_{n})$
$=$ $\frac{1}{|\mathcal{D}^{\mathsf{Va}}_{n}|}\sum_{d\in\mathcal{D}_{n}^{\mathsf{Va}}}\ell(\bm{\omega}_{n}^{(t),k};d),\bm{\omega}_{n}^{(t),k}\in\mathbb{R}^{|\bm{\omega}_{n}|}.$ (9)
We similarly define the training loss for institution $n$’s model at SGD iteration $k$ of global aggregation step $t$ as follows:
$\mathcal{L}_{n}^{\mathsf{Tr}}(\bm{\omega}_{n}^{(t),k})\triangleq$ $\mathcal{L}_{n}(\bm{\omega}_{n}^{(t),k};\mathcal{D}^{\mathsf{Tr}}_{n})$
$=$ $\frac{1}{|\mathcal{D}^{\mathsf{Tr}}_{n}|}\sum_{d\in\mathcal{D}_{n}^{\mathsf{Tr}}}\ell(\bm{\omega}_{n}^{(t),k};d),\bm{\omega}_{n}^{(t),k}\in\mathbb{R}^{|\bm{\omega}_{n}|}.$ (10)
We next derive two metrics called overfitting $O_{C}^{(t)}$ and generalization $G_{C}^{(t)}$ for each modality combination $C$ across the system, where $C\in\mathcal{P}(\mathcal{M})$, defined as
$O_{C}^{(t)}$ =¯$=\overline{\mathcal{L}}_{C}^{\mathsf{Va},(t)}-\overline{\mathcal{L}}_{C}^{\mathsf{Tr},(t)},$ (11)
$G_{C}^{(t)}$ =¯$=\overline{\mathcal{L}}_{C}^{\mathsf{Va},(t)},$ (12)
where ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Va},{(t)}}$ and ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Tr},{(t)}}$ are the average of the train and validation loss from institutions training on data with modality combination $C=\mathcal{M}_{n}$, obtained based on (IV-B) and (IV-B) as later formalized through PCW in Sec. IV-C. At each global aggregation step $t$, the above two metrics are used in defining another metric, called Distributed Overfitting to Generalization Ratio (DOGR) denoted by $\Gamma_{C}^{(t)}$, defined as follows:
$\Gamma_{C}^{(t)}=\frac{1}{\varphi^{(t)}}\times\frac{\left|\Delta G_{C}^{(t)}\right|}{\left(\Delta O_{C}^{(t)}\right)^{2}},$ (13)
where $\varphi^{(t)}=\sum_{C\in\mathcal{P}(\mathcal{M}_{n})}\frac{|\Delta G_{C}|}{2(\Delta O_{C})^{2}}$ is a normalization coefficient. Also, $\Delta O_{C}^{(t)}=O_{C}^{(t)}-O_{C}^{(t-1)}$ and $\Delta G_{C}^{(t)}=G_{C}^{(t)}-G_{C}^{(t-1)}$ represent the variations in overfitting and generalization, respectively. In (13), $\Delta O_{C}^{(t)}$ captures the change in the overfitting during the $t$-th global aggregation step for modality subset $C$ under the used step size. Also, $\Delta G_{C}^{(t)}$ captures the change in the validation loss during the latest global aggregation step for modality subset $C$, which measures the overall performance variation of the respective modality subset. Consequently, $\Gamma_{C}^{(t)}$ is a metric that takes smaller values when the generalization of the modality subset $C$ is decreasing or the model is becoming overfit during the global round and takes larger values otherwise. Thus, we can use $\Gamma_{C}^{(t)}$ to accelerate and slacken the learning-rate of different modality subsets according to their exhibited performance by increasing and dampening the unified SGD step-size during each local training round as discussed below.
We incorporate $\Gamma_{C}^{(t)}$ in the learning-rate of the local SGD iterations at each institution $n$. In particular, we modify the local SGD iterations given by (5) to the following update rule, which accelerates/slackens the learning-rate across different modality subsets of various institutions ($10\leq k\leq K-$ 1):
$\bm{\omega}^{(t),k+1}_{m,n}=\bm{\omega}^{(t),k}_{m,n}-\eta^{(t),k}\Gamma_{m}^{(t)}\widetilde{\nabla}_{m}\mathcal{L}_{n}(\bm{\omega}^{(t),k}_{n};\mathcal{B}_{n}^{(t),k}),\forall m\in\mathcal{M}_{n},$ (14)
where ~$\widetilde{\nabla}_{m}$ is the stochastic gradient with respect to the parameters of modality $m$’s encoder (note $\Gamma_{m}^{(t)}=\Gamma_{C}^{(t)}$ with $C=m$). The classifier update is also modified as follows:
$\bm{\varpi}^{(t),k+1}_{n}=\bm{\varpi}^{(t),k}_{n}-\eta^{(t),k}\hskip 0.85358pt\Gamma_{\mathcal{M}_{n}}^{(t)}\widetilde{\nabla}_{\bm{\varpi}}\mathcal{L}_{n}(\bm{\omega}^{(t),k}_{n};\mathcal{B}_{n}^{(t),k}),$ (15)
where $\Gamma_{\mathcal{M}_{n}}^{(t)}$ ($\Gamma_{\mathcal{M}_{n}}^{(t)}=\Gamma_{C}^{(t)}$ with $C=\mathcal{M}_{n}$) captures the DOGR for the subset of modalities at institution $n$ and ~$\widetilde{\nabla}_{\bm{\varpi}}$ is the stochastic gradient with respect to its classifier parameters.
Naïve implementation of DGB for longer local round lengths (i.e., $K$) in our system will result in biased ML models (due to the heterogeneity of local datasets) being aggregated to estimate the gradient blending weights. Thus, one of the conditions that needs to be met in gradient blending is that the dataset used to estimate the weights should follow the same distribution as the general dataset that all of the institutions possess, which corresponds to the dataset $\mathcal{D}$ in our system. Meanwhile, one of the general phenomena in FL is the non-iid-ness of the local datasets. This, contrary to the requirements of DGB, makes the weights estimated unreliable in the sense that institutions with different ML model biases will have the same contribution to DGB. This is why we propose PCW in order to control the contribution of each institution’s loss to the gradient weight estimation based on the distance between their data distribution $\mathcal{D}_{n}$ and $\mathcal{D}$. In doing so, we will define a dynamic metric that conceptualizes the distance between institutions data distributions, and manipulates the weight with which each $\mathcal{L}_{n}^{\mathsf{Tr},(t)}$ and $\mathcal{L}_{n}^{\mathsf{Va},(t)}$ is taken into effect.
Remark 1 (Learning Methodology, Unbalanced Number of Modalities, Robustness to Non-iid data and Data Scarcity).
The issue of varying convergence rates across modalities has also been observed in conventional centralized multi-modal ML [21]. To address this issue, [21] proposed a methodology, called gradient blending (GB), for centralized ML settings, which entails the use of a separate loss for each modality and a multi-modal model loss, weighting gradients non-uniformly when back-propagated over the ML model during training. As a result, since our methodology is inspired by [21], although it contains multiple nuances and differences compared to the method developed in [21], we call it Distributed Gradient Blending (DGB). In particular, our method differs significantly from the naïve extension of GB to the FL setting as done in [31]. In particular, existing methods [31, 21] use multiple sub-networks (i.e., one set of encoder and classifier for each modality), and thus, multiple losses to back-propagate during ML model training. In contrast, our work does not entail using multiple sub-networks; instead, we use only one classifier in each institution’s ML model alongside separate encoders dedicated to various modalities. We further use only one loss function and weigh its resulting gradient vector elements based on the encoder they belong to. This reduces the computation complexity of our method and the number of model parameters to train in each institution. Also, in addition to having a higher number of parameters, in the naïve implementations of gradient blending [21, 31], the methods require the train data available ($\mathcal{D}_{n}^{\mathsf{Tr}}$) to be split into two sections, one of which is used for the calculation of the overfitting to generalization ratio. This raises challenges when facing a scenario with scarce data, which is frequent in medical domains given the labor needed to annotate data. Our method, on the other hand, uses system-wide characteristics and values of the losses to calculate the DOGR which removes the necessity to have each of the train and validation data to be split into two at each institution, enabling more compatibility with data-scarce scenarios. More importantly, we address the side-effects of non-iid data on the performance of multi-modal FL through PCW, which makes DGB outperform the existing methodologies.
IV-C Proximity-Aware Client Weighting (PCW)
We propose PCW to tackle the problem of non-iid distribution of local datasets, which leads to the bias of locally trained models in multi-modal FL. PCW defines a measure to weigh the institutions’ local model losses (i.e., $\mathcal{L}_{n}^{\mathsf{Va},(t)}$ and $\mathcal{L}_{n}^{\mathsf{Tr},(t)}$ given in (IV-B) and (IV-B)) to obtain the normalized losses (i.e., ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Va},(t)}$ and ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Tr},(t)}$ used in (11) and (12), and subsequently (13). This measure is defined based on the quality of local datasets of institutions. In order to obtain this measure, we rely on a key observation: considering that the ML models at institutions are initialized with a unique global model at the beginning of each local training round, the difference in their local gradient trajectories is an indicator of the difference in their local data distributions. Furthermore, the difference/closeness between the local gradient trajectories and the gradient trajectory of the global model is an indicator of the similarity of local datasets to the global dataset (i.e., institutions with higher similarity of local gradient trajectory to that of the global model have a closer local data distribution to that of global dataset).
To realize PCW, we thus aim to compute the gradient trajectories at institutions and their similarities to that of global model. Therefore, we calculate the cumulative gradient of institution $n$’s model possessing modality combination $C=\mathcal{M}_{n}$ during the latest global aggregation step, denoted by $\Delta\bm{\omega}_{C,n}^{(t)}$, and then compute the trajectory of the global model gradient, denoted by $\Delta\bm{\omega}_{C,G}^{(t)}$. Mathematically, $\Delta\bm{\omega}_{C,n}^{(t)}$ and $\Delta\bm{\omega}_{C,G}^{(t)}$ are given by (note that $\bm{\omega}_{n}^{(t),0}=[\{\widehat{\bm{\omega}}_{m}^{(t)}\}_{m\in C}:\widehat{\bm{\varpi}}_{C}^{(t)}]$)
$\Delta\bm{\omega}_{C,n}^{(t)}$ $=\Big{(}\bm{\omega}_{n}^{(t),0}-\bm{\omega}_{n}^{(t),K}\Big{)}=\Big{(}\bm{\omega}_{n}^{(t),0}-\bm{\omega}_{n}^{(t),1}\Big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\Big{(}\bm{\omega}_{n}^{(t),1}-\bm{\omega}_{n}^{(t),2}\Big{)}+\cdots+\Big{(}\bm{\omega}_{n}^{(t),K-1}-\bm{\omega}_{n}^{(t),K}\Big{)}$
$=\sum_{j=0}^{K-1}\eta^{(t),j}\widetilde{\nabla}\mathcal{L}_{n}(\bm{\omega}^{(t),j}_{n};\mathcal{B}_{n}^{(t),j})$ (16)
and similarly
$\Delta\bm{\omega}_{C,G}^{(t)}=[\{\widehat{\bm{\omega}}^{(t)}\}_{m\in C}:\widehat{\bm{\varpi}}_{C}^{(t)}]-[\{\widehat{\bm{\omega}}^{(t+1)}\}_{m\in C}:\widehat{\bm{\varpi}}_{C}^{(t+1)}].$ (17)
Next, to compute the similarity between the local and global gradient trajectories, we adopt the inner product operator in defining a similarity/closeness metric as follows:
$\rho_{C,n}^{(t)}=\langle\Delta\bm{\omega}_{C,n}^{(t)},\Delta\bm{\omega}_{C,G}^{(t)}\rangle=\Delta\bm{\omega}_{C,n}^{(t)}\cdot\Delta\bm{\omega}_{C,G}^{(t)}.$ (18)
Considering (18), $\rho_{C,n}^{(t)}$ takes larger values when local gradient trajectory $\Delta\bm{\omega}_{C,n}^{(t)}$ exhibits a higher similarity to that of global gradient trajectory $\Delta\bm{\omega}_{C}^{(t)}$. [^5] In addition to inner product, we also experimented with other metrics of similarity between the vectors, such as the infimum of the difference in the parameters. Nevertheless, inner product exhibited the best result in our experiments, and thus the rest of the metrics are omitted for brevity. To have comparable values of $\rho_{C,n}^{(t)}$ across institutions $n\in\mathcal{N}$, we normalize the values obtained in (18) to obtain a normalized similarity metric as follows:
$\overline{\rho}_{C,n}^{(t)}={\rho_{C,n}^{(t)}}\Big{/}\Big{(}{\sum_{n\in\mathcal{N}}b_{C,n}\rho_{C,n}^{(t)}}\Big{)},$ (19)
where $b_{C,n}$ is a binary indicator that takes the value of $1$ when where the modalities at institution $n$ match the combination $C$ (i.e., $\mathcal{M}_{n}=C$); otherwise $b_{C,n}=0$.
Having the weights for each institution possessing subset $C$ of modalities, we then derive the average train and validation loss of the respective subset of modalities in the system at the end of each global aggregation round $t$, denoted by ¯$\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}$ and ¯$\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}$, respectively, as
¯$\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}=\frac{1}{\sum_{n\in\mathcal{N}}b_{C,n}}\sum_{n\in\mathcal{N}}\overline{\rho}_{C,n}^{(t)}b_{C,n}\mathcal{L}^{\mathsf{Tr},(t)}_{n},$ (20)
¯$\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}=\frac{1}{\sum_{n\in\mathcal{N}}b_{C,n}}\sum_{n\in\mathcal{N}}\overline{\rho}_{C,n}^{(t)}b_{C,n}\mathcal{L}^{\mathsf{Va},(t)}_{n}.$
Using the estimates in (20), we can substitute the values for ¯$\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}$ and ¯$\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}$ in (11) and (12) to calculate the overfitting and generalization, and consequently we can calculate the weights in (13). Compared to a naïve averaging of the training and validation losses, this new formulation of ¯$\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}$ and ¯$\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}$ enhances the performance by weighing losses from institutions with closer data distribution to $\mathcal{D}$ with higher values. This is an important step since the average losses being calculated in (11) and (12) will be used for the entire system in (13). Therefore, the institutions where data induces more bias should be given a smaller weight/importance when calculating the average loss, which will determine the DGB weights for all institutions. Our overall methodology for multi-modal FL comprising DGB and PCW is provided in Algorithm 1.
V Numerical Experiments
In this section, we first detail the datasets in Sec. V-A. Sec. V-B outlines the preprocessing steps undertaken for each data modality. We then review the baseline methods, describing their workings, result replication, and their relevance to our scenario in Sec. V-C. Sec. V-D describes the tuning of hyperparameters. Finally, simulation results are presented in Sec. V-E, complemented with discussions on their implications.
input : Clients $\mathcal{N}$, Datasets of clients $\{\mathcal{D}_{n}\}_{n\in\mathcal{N}}$, Number of global steps $T$, Number of local SGD iterations $K$ in each global step, Initial weights for gradient blending $\Gamma^{(0)}$
1 for global aggregation round $t=1$ to $T$ do
2       % On the Clients’ Side
3       for client $n=1$ to $|\mathcal{N}|$ in parallel do
4             if $t<2$ then
5                   Set the encoders’ learning rate coefficients based on $\Gamma^{(0)}$
6            else
7                  Calculate the encoder learning rate coefficients $\Gamma_{m},\forall m\in\mathcal{M}_{n}$ based on (13)
8            Perform $K$ rounds of local SGD based on (14) and (15).
9             Send the train loss $\mathcal{L}_{n}^{\mathsf{Tr}}(\bm{\omega}_{n}^{(t),K})$, validation loss $\mathcal{L}_{n}^{\mathsf{Va}}(\bm{\omega}_{n}^{(t),K})$, and the model weights $\bm{\omega}_{n}^{(t),K}$ to the server
10      % On the Server’s Side
11       for modality $m=1$ to $|\mathcal{M}|$ do
12             Aggregate the encoders with weight $\bm{\omega}_{m,n}^{(t),K}$ based on (7).
13      for classifier $c=1$ to $|\mathcal{P}(\mathcal{M})|$ do
14             Aggregate the classifiers with weight $\bm{\varpi}_{n}^{(t),K}$ based on (8).
15      for modality combination $C\in\mathcal{P}(\mathcal{M})$ do
16             calculate the weight $\overline{\rho}_{m,n}^{(t)}$
17             if $t\geq 2$ then
18                   calculate the average training and validation losses in global round $t$, ¯$\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}$ and ¯$\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}$ based on (20)
19
20
21      Send the calculated $\bm{\omega}_{m,n}^{(t+1),0}$, $\bm{\varpi}_{C}^{(t+1),0}$, {¯$\{\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}\}_{C\in\mathcal{P}(\mathcal{M})}$, and {¯$\{\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}\}_{C\in\mathcal{P}(\mathcal{M})}$ back to the clients.
Algorithm 1 Multi-modal FL over Institutions with Unbalanced Modalities
V-A Dataset and Data Distribution
We use datasets from The Cancer Genome Atlas program (TCGA) for three cancer types, namely, Breast Invasive Carcinoma (BRCA), Lung Squamous Cell Carcinoma (LUSC), and Liver Hepatocellular Carcinoma (LIHC), to make up for the heterogeneity of local datasets across clients. Each dataset has three modalities: mRNA sequence, histopathological images, and clinical information. Only the patient data that have all three modalities to make the results comparable to centralized ML as conducted in [9] and the FL settings described in [31, 32]. Nevertheless, in our setting, as will be described later, the data is distributed across the system such that each institution has access to only a fraction of the data modalities. The cancer stage distribution for the three datasets is described in Table I.
TABLE I: Data distribution across cancer stages in the datasets used.
Stage Type
BRCA
LUSC
LIHC
Stage 1
190
66
51
Stage 2
55
45
39
All experiments were conducted on a server with NVIDIA’s Tesla A-100 GPUs. For software implementation, we used PyTorch [35] module using Python programming language. Also, our codes are available at https://github.com/KasraBorazjani/dgb-pcw-fl.git.
V-B Preprocessing Phase
As dealing with data in raw format is highly challenging, especially in the case of histopathological images (see Fig. 4) where data is of large size, we follow the method of [36] to preprocess data. In this procedure, the nucleus segmentation is done using the unsupervised method of [37]. We then extract $10$ cell-level features, namely, (i) nucleus area, (ii) major axis length, (iii) minor axis length, (iv) major to minor axis ratio, (v) mean distance to the nearby nuclei, (vi) maximum distance to the nearby nuclei, (vii) minimum distance to the nearby nuclei, and the average value of each of the (viii) red, (ix) green, and (x) blue channels in the surrounding area of the nuclei. These will constitute our cell-level features. To make a patient-level or slide-level representation for each patient, we take each of cell-level features and create a $10$-bin histogram for it, alongside the five metrics of mean, standard deviation, skewness, kurtosis, and entropy. Taking each bin of the histogram and the aforementioned metrics as a patient-level feature will yield $150$ features in total which we use for describing a single histopathological image (WSI) for each patient. The rest of data modalities (i.e., mRNA sequence and clinical information) are fed to the ML model in raw formats.
Fig. 4: A sample of the histopathological image, containing dense pixels.
V-C Baseline Methods
We compare the results of our method with three baselines:
  1. 1.
    Conventional Multi-Modal FL (CM-FL): Following the model in Fig. 1, institutions have heterogeneous modalities and local ML models are trained using the same learning-rate and aggregated using the FedAvg algorithm [32].
  2. 2.
    Multi-Modal FL with Hierarchical Gradient Blending (HGB-FL): We consider [31] to test our method versus another implementation of gradient bending for FL. This approach splits the train data at each institution to calculate the gradient blending weights and tunes the weights with which the ML models will be aggregated at the server.
  3. 3.
    CM-FL + DGB: The institutions will possess heterogeneous modalities and their ML models will be trained using DGB without the use of PCW.
  4. 4.
    UM-FL: The upper bound FL performance is achieved when all institutions possess all of data modalities, removing the barrier of different convergence rates across modalities. This baseline is only added to provide an estimate of the achievable performance.
We consider the first baseline to measure the system’s performance under the conventional methods (i.e., FedAvg aggregation and local training using SGD), which is the naïve extension of FL to multi-modal setting as described in [32]. We will demonstrate that this approach will perform poorly over unbalanced modalities across the institutions. The second baseline [31] implements a federated version of gradient blending introduced in [21], which is not tailored to address the problem of uneven number of modalities across institutions. This baseline also relies on tuning the global model aggregations weights, which is different from our method. Further, our method removes the necessity to split the training data at each institution, as done in [21], to calculate the gradient blending weights. This is reflected in our formulation (20) where we aggregate all of the loss values needed to calculate the DGB parameters based on the modality combination of the institutions to get a distributed calculation of these parameters. Avoiding splitting the dataset in data-scarce scenarios such as ours, and generally, most medical applications, can (later shown in Table III) improve the outcome of the federated learning. We further complement our method with PCW, which as will be shown later will give us performance edge. In the third baseline, we aim to measure the base performance of our proposed DGB and demonstrate the performance gain we get by employing the method. Although we use this method as a baseline, we will later show that this method outperforms the state-of-the-art, e.g., [32, 31], due to consideration of different convergence speeds across various modalities.
V-D Hyperparamters Tuning
The major hyperparameters tuned in our simulations are the number of layers for the mRNA modality’s encoder depth, the initial learning-rate and its decay-rate. For the mRNA modality’s encoder’s depth, we conducted three experiments with three neural network architectures with multiple layer depths which are listed in Table II. For the initial learning-rate, we chose five values to start the ML training with $\{10^{-2},10^{-3},10^{-4},10^{-5},10^{-6}\}$. The decay-rates for the learning-rate were searched between the values $\{0.99,0.95,0.90,0.85\}$. The best performance for different mixtures of these parameters was observed for the first $50$ global FL steps and the medium depth, $10^{-4}$ initial learning rate, and $0.99$ decay rate over each global round were chosen. We used $K=20$ local SGD iterations on each institution’s ML model. Also, dividing the datasets mentioned in Sec. V-A, we simulated $9$ institutions with different subsets of modalities.
TABLE II: Different architectures for the mRNA encoder.
Layer Type
Model Depth
Small
Medium
Large
Input Layer
FC (20531, ReLU)
Layer 1
FC (4096, ReLU)
FC (8192, ReLU)
FC (16384, ReLU)
Layer 2
FC (2048, ReLU)
FC (4096, ReLU)
FC (8192, ReLU)
Layer 3
FC (512, ReLU)
FC (2048, ReLU)
FC (4096, ReLU)
Layer 4
FC (128, ReLU)
FC (512, ReLU)
FC (2048, ReLU)
Layer 5
FC (64, ReLU)
FC (128, ReLU)
FC (512, ReLU)
Layer 6
FC (32, ReLU)
FC (64, ReLU)
FC (128, ReLU)
Layer 7
FC (32, ReLU)
FC (64, ReLU)
Layer 8
FC (32, ReLU)
Output Layer
FC (2, LogSoftmax)
V-E Results and Discussions
We next present our simulation results and their implications.
V-E1 ML training under different learning methods
To test the performance of our method against the baselines, we devised an experiment trying to simulate the methods under the same circumstances. In particular, all of the ML models start from the same initial point/model in order to avoid having random initial weights affect the performance of the models. We took the datasets mentioned in Sec. V-A and created different subsets from each of them to account for the multiple institutions that we have in our network. We acquired the results gathered in Table III for the evolution of performance of each method over time. As can be seen, CM-FL + DGB already surpasses the CM-FL method. However, as the training proceeds, using our entire methodology with the addition of PCW, we can see a better performance. This reveals that the client weighting method used in PCW, although it can slow down the training at the beginning, can effectively circumvent the local ML model bias, especially as ML training proceeds. The table further shows that, due to the unique challenges introduced by the unbalanced distribution of data modalities and labels, baselines that are not designed to handle such non-uniformaties (CM-FL, HGB-FL) fail to train the model effectively.
TABLE III: Accuracy of different methods throughout the global aggregations.
Global Round Method
CM-FL
HGB-FL
CM-FL + DGB
Our Method
UM-FL (Upperbound)
1
$\simeq$ 61.80
$\simeq$ 48.72
$\simeq$ 63.49
$\simeq$ 39.32
$\simeq$ 51.34
100
$\simeq$ 62.92
$\simeq$ 67.52
$\simeq$ 63.34
$\simeq$ 61.80
$\simeq$ 50.89
200
$\simeq$ 59.55
$\simeq$ 65.81
$\simeq$ 63.64
$\simeq$ 64.04
$\simeq$ 51.34
300
$\simeq$ 62.92
$\simeq$ 63.25
$\simeq$ 64.24
$\simeq$ 68.54
$\simeq$ 66.52
400
$\simeq$ 61.80
$\simeq$ 62.39
$\simeq$ 66.43
$\simeq$ 71.91
$\simeq$ 79.46
500
$\simeq$ 64.04
$\simeq$ 64.10
$\simeq$ 67.13
$\simeq$ 74.16
$\simeq$ 79.02
600
$\simeq$ 61.80
$\simeq$ 62.39
$\simeq$ 69.21
$\simeq$ 74.16
$\simeq$ 79.02
V-E2 ML training under different levels of non-uniformity across the modalities
Since we are among the first to study an FL system with an unbalanced number of modalities across clients/institutions, we study the impact of varying the non-uniformity of the number of modalities across institutions. We compare the performance of our method against the baselines under different levels of non-uniformity to show the robustness of our method against the distribution of modalities across institutions. Since we have three modalities, we can have either uni-modal clients, bi-modal clients, or tri-modal clients. We conduct our comparisons in three settings where most institutions possess the dominant number of modalities:
  1. 1.
    Uni-modal dominant: $2/3$ of institutions have uni-modal data, $1/6$ of institutions have bi-modal data, and $1/6$ of institutions have tri-modal data.
  2. 2.
    Bi-modal dominant: $1/6$ of institutions have uni-modal data, $2/3$ of institutions have bi-modal data, and $1/6$ of institutions have tri-modal data.
  3. 3.
    Tri-modal dominant: $1/6$ of institutions have uni-modal data, $1/6$ of institutions have bi-modal data, and $2/3$ of institutions have tri-modal data.
We compare the performance (in terms of prediction accuracy) of our method against the baselines in Table IV. It can be seen that our method outperforms the other baselines by a notable margin. Also, reducing the number of modalities per client (i.e., having all tri-modal clients vs. having all uni-modal clients) leaves a marginal impact on the performance of our method (i.e., $1.33\%$) as compared to baselines, being $1.76\%$ and $2.13\%$. This implies the robustness of our method against the level of heterogeneity of the modalities across institutions.
TABLE IV: Global model accuracy for each dominant modality scenario.
Method Dominant Modality
Uni-Modal
Bi-Modal
Tri-Modal
CM-FL
$\simeq$ 60.12
$\simeq$ 60.56
$\simeq$ 62.25
CM-FL + DGB
$\simeq$ 69.44
$\simeq$ 70.13
$\simeq$ 71.20
Our Method
$\simeq$ 74.30
$\simeq$ 74.83
$\simeq$ 75.63
V-E3 ML training under different number of modalities involved
In addition to heterogeneity of number of modalities across the system, the quality of data of each modality is different across institutions (i.e., some modalities may contribute more to the overall performance of the ML model as compared to the rest). To investigate this, we train the ML models and gather the accuracy of the global model under availability of different subsets of modalities across institutions in Table V. The results show that our method outperforms the baselines under various data qualities. Also, the clinical data shows the lowest quality data as it leads to the lowest performance when combined with other modalities, while the mRNA modality has the highest quality data. This observation can further open the door to research on multi-modal FL in the presence of different quality data and attention mechanisms, which can put non-uniform emphasis on different types of data.
TABLE V: Accuracy of different methods under different subsets of modalities.
Method Modality Subset
mRNA+Image + Clinical
mRNA + Image
mRNA + Clinical
Image + Clinical
CM-FL
$\simeq$ 61.80
$\simeq$ 61.25
$\simeq$ 55.32
$\simeq$ 50.43
CM-FL + DGB
$\simeq$ 69.21
$\simeq$ 65.23
$\simeq$ 55.81
$\simeq$ 54.43
Our Method
$\simeq$ 74.16
$\simeq$ 68.42
$\simeq$ 58.12
$\simeq$ 55.84
VI Conclusion and Future Works
In this work, we developed the system model for multi-modal FL over institutions with non-iid data and unbalanced number of modalities. To address disparate convergence rates across the various modalities involved in the system, we proposed DGB, weighing the gradients of different modalities in a non-uniform fashion. Further, to address the impact of non-iid data and the local model bias on the performance of DGB, we proposed PCW which takes into account the quality of data at institutions for refining the system-wide loss estimates fed to DGB. We demonstrated that our proposed methodology can outperform the state-of-the-art methods. Our proposed DGB and PCW methods are generic and can be used in other tasks such as Alzheimer’s disease prediction with data modalities of CT and PET scans, the exploration of which is left as future work. Other future directions include using generative adversarial networks (GANs) to replicate the missing modalities at each institution. Also, introducing distributed modality-aware attention mechanisms is a promising research direction.